Two masses collide on a frictionless horizontal floor and in perfectly inelastic collision. Mass 1 is 4 times that of mass 2. Velocity of mass 1 = 10 m/s to the right while the velocity of mass 2 = 20 m/s to the left. What is the velocity and direction of the resulting combined mass?
1
Expert's answer
2021-08-30T15:05:35-0400

Let's choose the rightwards as the positive direction. Then, we can find the velocity and direction of the resulting combined mass from the law of conservation of momentum:


"m_1v_1-m_2v_2=(m_1+m_2)v_c,""v_c=\dfrac{m_1v_1-m_2v_2}{m_1+m_2}."

Since "m_1=4m_2" we get:


"v_c=\dfrac{4m_2v_1-m_2v_2}{4m_2+m_2}=\dfrac{m_2(4v_1-v_2)}{5m_2}=\dfrac{4v_1-v_2}{5},""v_c=\dfrac{4\cdot10\ \dfrac{m}{s}-20\ \dfrac{m}{s}}{5}=4\ \dfrac{m}{s}."

The sign plus means that the velocity of the resulting combined mass directed to the right.